The following information relates to Riggs Corp.’s purchase of equipment on 15 June 2017:

Invoice price

$

448,000

Repairs required in the first month of use

2,400

Shipping costs

2,930

Installation

5,000

Testing

10,000


The equipment was installed and tested during the week of 22 June 2017. Riggs paid the invoice price on 1 July 2017. The equipment was ready for use on 30 June and put into production on 3 July 2017. Riggs uses straight-line depreciation for the company’s equipment and expects to use the asset for six years. Component parts are not significant and need not be recognized and depreciated separately. The estimated residual value is zero. The company’s fiscal year-end is 31 December.

Required:
1. Prove that the book value which Riggs will report for the equipment purchase is $465,930. Show your work.


2. Compute depreciation expense for 2017, using the straight-line method, under each of the following assumptions: (Round your answers to nearest whole dollars.)

  1. Exact, to the closest month
  1. Full first-year convention

  1. Half-year convention

3. Prepare the journal entry to record depreciation expense, assuming Riggs uses the straight-line method and the half-year convention.

General Journal

Date

Account Titles and Explanation

Debit

Credit
Answer #1

1)

Invoice price 448000
Shipping costs 2930
Installation 5000
Testing 10000
Purchase cost 465930

2)Depreciation expense = [Cost -Salvage value if any]/Useful life

                            = [465930 - 0 ]/ 6

                             = $ 77655

a)Exact ,closest to month :Depreciation will be charged from the date ,asset is ready for use that is from 30 June .

Depreciation for 2017 (30June -31Dec = 6 months ) = 77655*6/12 = 38827.5 (rounded to 38828)

b)Full first year convention :Irrespective of month of acquisition ,depreciation is allowable for complete /whole year = $ 77655

c)

Half year convention :This method assumes asset is acquired in the middle of year and depreciation is charged for the half year accordingly .

Depreciation expense= 77655*6/12 = 38828

3)

Date Account ttitle Debit credit
31December 2017 Depreciation expense 38828
Accumulated depreciation 38828